OpenStudy (anonymous):

write an equation in slope intercept form of the line that passes through thhe point (-6,-5),(1,4)

OpenStudy (anonymous):

find the slope: rise over run then plug it into y= mx+ b form m= slope solve for b

OpenStudy (anonymous):

m= y2 -y1 over x2- x1

OpenStudy (anonymous):

ineed to know how to doo it step by step idnt knw hw

OpenStudy (anonymous):

ok use your points... and plug them in...\[4--5 \over 1- -6\]

OpenStudy (anonymous):

solve that... and tell me what you got.

OpenStudy (anonymous):

ok i got -1 over -5

OpenStudy (anonymous):

ok i got 9/7.... if you plug it into a calc. 4-(-5) = 9 and 1-(-6) i got 7

OpenStudy (anonymous):

o sorry i did it wrong u right ok what next?

OpenStudy (anonymous):

so that is M in your equation...

OpenStudy (anonymous):

so we are going to use the points (1,4) 4= 9/7(1) + B now solve for b.

OpenStudy (anonymous):

b=0?

OpenStudy (anonymous):

is b=0 right man

OpenStudy (anonymous):

ok you do (9/7) x 1 and then you bring that to the side where 4 is... by subtracting and that gives you b.